What is the specific gravity of brandy 41% the specific gravity of gin 38% and the specific gravity of tequila 50%
Answer #1

Brandy (41%), gin (38%) and tequila (50%) are all forms of alcoholic beverages with the mentioned percentage composition of alcohol and water.

As we know, density = mass / volume

Therefore, the density of the mixture of water and alcohol would be = (M1+M2) / (M1/D1+M2/D2)

where, M1 = mass of alcohol

M2 = mass of water

D1 = density of alcohol

D2 = density of water

Dividing the overall density of the mixture by the density of water will give the specific gravity of the mixture. As the density of water is approximately equal to unity, so the values of the density of mixture would be the same as the specific gravity of the mixture.

Calculation of specific gravity of brandy

Density of brandy (41%) = (M1+M2) / (M1/D1+M2/D2)

= (41 + 59) / (41 / 0.789 + 59 / 1)

= 100 / (51.96 + 59)

= 100 / 110.96

= 0.901  

So, the specific gravity of brandy (41%) = 0.901

Calculation of specific gravity of gin

Density of gin (38%) = (M1+M2) / (M1/D1+M2/D2)

= (38 + 62) / (38 / 0.789 + 62/ 1)

= 100 / (48.16 + 62)

= 100 / 110.16

= 0.908

So, the specific gravity of brandy (41%) = 0.908

Calculation of specific gravity of tequila

Density of brandy (50%) = (M1+M2) / (M1/D1+M2/D2)

= (50 + 50) / (50 / 0.789 + 50 / 1)

= 100 / (63.37 + 50)

= 100 / 113.37

= 0.882  

So, the specific gravity of brandy (41%) = 0.882